OpenStudy (anonymous):

Renee needs to determine the equation of a line passing through the point (-1, 7) that is parellel to y=-3x-4. What is the proper equation?

OpenStudy (btaylor):

use \[(y-y_{1}) = m(x-x_{1})\] Since the lines are parallel, the slopes are the same. So you get:\[(y-7)=-3(x+1)\] Simplified: \[y=-3x-3+7 = -3x+4\]

OpenStudy (anonymous):

thanks....i get confused by the formula sometimes....